From GlueXWiki
Jump to: navigation, search

January 9, 2014 FDC meeting


  1. Installation [1],[2],[3],[4] (all)
  2. Engineering [5], [6], [7] (Bill)
  3. Electronics (Chris, Nick)
  4. Mini-DAQ [8] (Beni)
  5. Other


Participants: Eugene, Fernando, Dave, Nick, Chris, Simon, Bill, Beni, and Lubomir.


- One quadrant (top, left looking downstream) has been completely cabled at the detector side (see attached pictures), main work done by Mike Klopf (FEL) with the help of Shukui Zhan (FEL) and Lubomir. Now Nick, Robert (FEL) and Shukui are working on the connections at the electronics side. As of today Casey and Tina started cabling the right top side. Cabling procedure: starting with first package per quadrant, each 11 cables are grouped, total of three groups naturally following the pre-amp positions, and bundled together starting at the end of the mesh. In total there are 12 bundles that are then placed on the top cable tray and then the end is re-bundled according to the rack/crate positions.

- In between, while the cables were on the floor, Anatoly tested all the cables at the electronics end with a scope using the translator board which allows to test only the first 8 and the last 8 channels from the 24 in the connector. There were four bad channels (small signals) in packages 3 and 4, most likely conductive tape issues. There's one card with all the channels showing a long shoulder in the signal, we suspect the pre-amp.

- Since Friday last weak we switched the gas from Nitrogen to Ar/CO2 90/10. All the HV channels were OK till Tuesday evening when the same channels (that showed problems before Holidays) starting tripping from time to time. Right now one of the HV channels is really problematic tripping at >1400V (1600V is the nominal). Another four channels are tripping once per 8-24 hours. Certainly the trips are related to the gas. On Monday we expect to get the nominal gas mixture, 40/60.


- Bill showed drawings (attached above) with the CDC bearings and attachment that require some space in between the two halves of the mesh. We wanted to put there three HV cables and a ground cable. Bill will get samples of these cable and look in place if they will fit there.

- Bill is trying to figure out the problem with the mass flow controller and contacted the Brooks company. The flow is about a factor of 2 lower than the nominal, but we are not sure how reliable the flow controller is.

- No results yet from the last survey done first week of December.

- Cathode scan: Bill will contact the company, expect a resolution of 100 microns which should be good enough because of the many measurements. Assumption is it will cost about few $k, Eugene agreed. The idea is to fit the scan results with a model with some free parameters that Simon will include in his alignment procedure.


- Fernando, Bill, Nick and Chris discussed how to do the "HV return" connection to ground (one per HV connector): will be connected to the ground plate (see Bill's pictures above). There will be also one ground cable connected to one of the two ground plates. The "HV shield" (one per HV connector) should be connected to one shield cable coming from a wire frame. All the other shield (green) cables that are not used (10 per package) should be insulated.


- At the meeting Beni explained he's very close to get the mini-DAQ working, and on the next day after the meeting Beni had already his first results. The trigger is generated by the software. There are two codes, one for calculating the pedestals and their widths, which is very fast. The other code analyses the events in the ROC and takes only those channels/events that had signals (within the time window) above a threshold. The statistics you can accumulate will depend on the time it takes to analyze the data in the ROC.